OpenStudy (anonymous):

Which options could be the next step in simplifying this expression? (18 + 23) + 7 Choose exactly two answers that are correct. A. 25 + 30 (use the distributive property) B. 30 + 25 (use the commutative property) C. 18 + 30 (use the associative property) D. 41 + 7 (use order of operations)

OpenStudy (anonymous):

Which property does the statement illustrate? Column A Column B 1. (p + 3) + q = p + (3 + q) 2. 8 + (9 + 4) = (8 + 9) + 4 3. 4(x – 2) = 4x – 8 4. 7(3 + 5) = 7 • 3 + 7 • 5 A. associative property of addition B. distributive property

OpenStudy (anonymous):

Which property does the statement illustrate? Column A Column B 1. (p + 3) + q = p + (3 + q) 2. 8 + (9 + 4) = (8 + 9) + 4 3. 4(x – 2) = 4x – 8 4. 7(3 + 5) = 7 • 3 + 7 • 5 A. associative property of addition B. distributive property